# Quantum Learning and Estimation for Distribution Networks and Energy Communities Coordination

### v1

@@ II-C Incentive-Response Mapping Integrated Gradient Descent for DNs-ECs Coordination
Under limited availability of information, DNs can only access the aggregated energy usage of each EC, while the detailed internal data and operational states of ECs remain unobservable. Therefore, the bilevel DNs-ECs coordination problem cannot be solved directly. To address this challenge, we propose to learn a direct mapping that captures the incentive-response behavior of ECs, as in (2).
$\hat{P}^{\mathrm{EX,EC}}_{i}=\varphi_{i}(\pi^{\mathrm{EC}};\varTheta)\sim\phi_{i}(\pi^{\mathrm{EC}}),$ (2)
@@
where the parameters $\varTheta$ can be optimized by minimizing a loss function (e.g., mean square error, MSE):
$\varTheta^{*}={\arg\min}_{\varTheta}\ \frac{1}{T}\sum_{t=1}^{T}\left(P^{\mathrm{EX,EC}}_{t}-\hat{P}^{\mathrm{EX,EC}}_{t}\right)^{2}.$ (3)
@@
Then, by using the mapping model in (2) to replace the unknown response function in (1h), the original bilevel DNs-ECs coordination problem is reformulated as a single-level optimization problem, yielding solutions equivalent to those of the original bilevel formulation.
With the differentiability of the mapping model (2), the integrated DNs Optimization model is differentiable and solvable via gradient-based methods. With respect to the decision variables $z$ and $v_{\alpha}$ of DNs optimization, the gradient descent update rule can be expressed as:
$z^{k+1}=z^{z}-\eta\frac{\partial C}{\partial z^{k}},\ \ v_{\alpha}^{k+1}=v_{\alpha}^{k}-\eta\frac{\partial C}{\partial v_{\alpha}^{k}},$ (4a)
$\frac{\partial C}{\partial z}=\sum_{s=1}^{S}\gamma_{s}\frac{\partial f(z,\xi_{s})}{\partial z}+\lambda\frac{1}{1-\alpha}\sum_{s=1}^{N}\gamma_{s}\frac{\partial f(z,\xi_{s})}{\partial z}I(f(z,\xi_{s})>v_{\alpha}),$ (4b)
@@
$\frac{\partial C}{\partial v_{\alpha}}=\lambda\left(1-\frac{1}{1-\alpha}\sum_{i=1}^{N}\gamma_{s}I(f(z,\xi_{s})>v_{\alpha})\right),$ (4c)
where $I(f(z,\xi_{s})>v_{\alpha})=1$ if $f(z,\xi_{s})>v_{\alpha}$ else 0.
@@
    The response behavior of ECs should be accurately estimated in (2). However, this task constitutes a complex long-horizon temporal sequence modeling problem, which poses significant challenges regarding both accuracy and generalizability.
@@
    To compute (4), numerous scenarios are required to achieve accurate risk quantification, which leads to a heavy computational burden.
In the next section, inspired by the recent advances in quantum computing, we propose a quantum learning and estimation approach to address the above two challenges, thus facilitating efficient DNs-ECs coordination.
@@ III Methodology of
In this section, we introduce the proposed quantum learning and estimation approach. Specifically, a Q-TCN-LSTM model is proposed to learn the incentive-response characteristics of ECs. Subsequently, a quantum estimation method is developed to accelerate the gradient computation across numerous scenarios.
III-A Structure of Q-TCN-LSTM
@@ III-A1 Variational Quantum Circuits
where $\left|\psi_{0}\right\rangle_{n}$ represents the initial quantum state of $n$ qubits; $\left|i\right\rangle_{n}$ is the $i$-th basis state in the $2^{n}$-dimensional Hilbert space, associated with amplitude $\alpha_{i}$ and a measurement probability of $\alpha_{i}^{2}$, which satisfies $\sum_{i=0}^{2^{n}-1}\alpha_{i}^{2}=1$. At this stage, VQC encodes exponentially many states via quantum superposition, thereby offering computational advantages in high-dimensional problems.
Secondly, the parameterized variational circuit block $V(\boldsymbol{\theta})$ applies a sequence of complex unitary transformations on $\left|\psi_{0}\right\rangle_{n}$. Single-bit gates (e.g., Pauli-X gate, rotation gate) and multi-bit gates (e.g., controlled-NOT gate, controlled-rotation gate) are used to manipulate the quantum state, collectively inducing quantum entanglement and further exploiting the superposition across qubits. This process enables the circuit to represent intricate nonlinear dependencies inherent in the input data. Usually, the variational circuit is composed of $L$ layers, each consisting of a set of parameterized or fixed single-qubit gates and entangling gates, forming the overall unitary:
@@
where $V^{\mathrm{ent}}_{(l)}$ is the entangling gate at layer $l$, $V^{\mathrm{rot}}_{(l)}(\boldsymbol{\theta}_{(l)})=\bigotimes_{i=1}^{n}R_{(x,y,z)}(\theta_{(l),i})$ is the rotation gate at layer $l$. With the variational circuit, the quantum state is transformed as:
@@
where $\left\langle\psi(\boldsymbol{\theta})\right|$ is the conjugate transpose of $\left|\psi(\boldsymbol{\theta})\right\rangle$. Equation (8) denotes the expectation value of the observable Hermitian operator $\mathcal{O}$, resulting in classical scalar values for post analysis.
The measurement output $\left\langle\mathcal{O}\right\rangle_{\boldsymbol{\theta}}$ represents a highly nonlinear transformation of the input, implicitly capturing complex correlations encoded by the VQC. The parameters $\boldsymbol{\theta}$ are optimized via classical optimization algorithms (e.g., Adam) to minimize a predefined loss function. Leveraging quantum parallelism and entanglement, VQCs can model complex patterns with significantly fewer trainable parameters compared to classical neural networks.
@@ III-A2 Structure of Q-TCN-LSTM
In response to time-varying electricity price signals, ECs exhibit both short-term (spanning several hours) response characteristics, and long-term (over the scheduling horizon) globally optimal decision-making. Therefore, a coupling relationship exists between the response of ECs and price incentives in both the short and long term. To capture this coupling, we propose a quantum-enhanced hybrid architecture, Q-TCN-LSTM, which integrates the advantages of temporal convolutional networks (TCNs) and long short-term memory (LSTM) networks. Specifically, TCNs can efficiently extract short-term features from time series data through causal and dilated convolution mechanisms. Meanwhile, LSTM can store long-term information in their cell states and selectively remember or forget information through gating mechanisms, effectively capturing long-term dependencies in time series. Fully connected (FC) layers then map the extracted features to the final output. The overall structure of the proposed Q-TCN-LSTM model is shown in Fig. 3.
@@ III-A3 Q-TCN for Short-Term Temporal Feature Extraction
Based on the classical TCN, we integrate VQCs into the causal convolutions to construct a quantum TCN (Q-TCN) model. For instance, a 4-qubit VQC for Q-TCN is illustrated in Fig. 4.
@@
The proposed Q-TCN model efficiently captures temporal correlations and complex nonlinear relationships across multiple time steps by embedding classical time-series data into a quantum feature space, leveraging quantum superposition and entanglement. Specifically, Q-TCN utilizes a temporal rotational encoding scheme, where each input value $x_{t}$ is mapped to the angle of a rotation gate. This encoding preserves local temporal patterns while mitigating potential information loss commonly associated with sparse sampling in classical TCNs. Then, a parameterized variational circuit with skip-connected entanglement gates performs entanglement across temporally aligned qubits, constructing temporal dependencies without deep stacking or activations in classical TCNs. Finally, quantum measurements produce output features that capture interactions across multiple time steps. Quantum properties such as superposition and entanglement enhance Q-TCN’s expressive power with significantly fewer trainable parameters, effectively reducing model redundancy.
III-A4 Quantum LSTM for Long-Term Temporal Feature Extraction
Based on the classical LSTM architecture, we construct a quantum LSTM (Q-LSTM) model by replacing the gate mechanisms with VQCs, as illustrated in Fig. 5.
Figure 5: Structure of Q-LSTM cell.
@@ III-A4 Quantum LSTM for Long-Term Temporal Feature Extraction
where $\sigma$ and $\tanh$ blocks represent the sigmoid and the hyperbolic tangent activation function, respectively. $x_{t}/h_{t}/c_{t}/y_{t}$ are the input, hidden state, cell state, and output at time $t$. $\otimes$ and $\oplus$ represents element-wise multiplication and addition. []delimited-[][$\cdot$][ ⋅ ] represents the vector concatenation, $f_{t}$, $i_{t}$, $o_{t}$ represent the forget gate, input gate, and output gate in the LSTM model, respectively.
III-B Quantum Amplitude Estimation
In (1a) and (4), a large number of scenarios are required for precise risk quantification, which introduces considerable computational complexity. Classical approaches typically rely on MC simulation for estimation, which has been verified to require $\mathcal{O}(1/\varepsilon^{2})$ samples to achieve an estimation error of $\varepsilon$. This leads to substantial computational overhead. Alternatively, quantum amplitude estimation, leveraging quantum entanglement and amplitude amplification, can theoretically achieve the same precision level $\varepsilon$ with only $\mathcal{O}(1/\varepsilon)$ queries [27], offering a theoretical quadratic speedup compared to MC. Next, we propose a quantum estimation method for efficient computation of (4). Take the first term in (4) (i.e., $\sum_{s=1}^{S}\gamma_{s}\frac{\partial f(x,\xi_{s})}{\partial x}$) as an example. We rewrite it as $\sum_{s=1}^{S}p(\xi_{s})\hat{f}(\xi_{s})$ and suppose $\xi$ follows distribution $\mathcal{X}$. The QAE procedure generally involves three steps:
Step 1-Quantum State Preparation: $\mathcal{X}$ is divided into $N=2^{n}$ intervals, and is embedded into a quantum state:
@@ III-B Quantum Amplitude Estimation
where $|i\rangle_{n}$ represents a basis state of $n$ qubits, $p_{i}$ is the probability of state $|i\rangle_{n}$ with $\sum_{i=0}^{N-1}p_{i}=1$. Common strategies for quantum state preparation include angle encoding, amplitude encoding, and basis encoding. In the uniform case where $p_{i}=1/N\ \forall i\in[0,\dots,N-1]$, (10) can be efficiently prepared using Hadamard gates.
Step 2-Quantum State Rotation: Assuming a function $\hat{f}:\{0,1,...,2^{n}-1\}\rightarrow[0,1]$, a controlled rotation operator $\mathcal{F}$ is constructed such that
$\mathcal{F}|i\rangle_{n}|0\rangle=\sqrt{1-\hat{f}(i)}|i\rangle_{n}|0\rangle+\sqrt{\hat{f}(i)}|i\rangle_{n}|1\rangle.$ (11)
@@
$\mathcal{F}|\psi\rangle_{n}|0\rangle=\sum_{i=0}^{N-1}\sqrt{(1-\hat{f}(i))p_{i}}|i\rangle_{n}|0\rangle+\sum_{i=0}^{N-1}\sqrt{\hat{f}(i)p_{i}}|i\rangle_{n}|1\rangle.$ (12)
@@
In the implementation of $\mathcal{F}$, the auxiliary qubit undergoes a rotation $R_{y}(\theta)$ with $\theta=2\arcsin(\sqrt{\hat{f}(i)})$. In our case, $\hat{f}(\xi)$ is a linear function of $\xi$, i.e., $\hat{f}(\xi)=a\xi+b$. Thus, $\theta=2\arcsin(\sqrt{a\xi+b})$, which is nonlinear with respect to $\xi$. Depending on the trade-off between circuit complexity and estimation accuracy, we propose two potential implementations of $\mathcal{F}$.
@@
The first circuit implementation, referred to as Circuit-1, employs analytical approximations of the nonlinear rotation and emphasizes low circuit complexity, as illustrated in Fig. 6. Assuming $x$ is sufficiently small, we have $\arcsin(\sqrt{x})\approx cx+d$, where $c$ and $d$ can be obtained via the least squares method. To conform to this approximation regime, a linear transformation can be applied to the function $\hat{f}(\xi)$, such that $\theta\approx\hat{a}\xi+\hat{b}$. Under this condition, the controlled rotation operator $\mathcal{F}$ can be efficiently realized using a series of parameterized controlled-RY gates, with low implementation complexity.
Figure 6: Circuit-1 for approximated implementation of $\mathcal{F}$.
@@
The second circuit implementation, referred to as Circuit-2, computes the exact nonlinear rotation angle corresponding to each quantum basis state. To this end, we design a rotation-angle-adaptive circuit architecture. In this scheme, $n$ original qubits are initialized in $|0\rangle_{n}$, $n$ reversal qubits in $|1\rangle_{n}$, and an auxiliary qubit in $|0\rangle$. The reversal qubits are kept in bitwise negated states of the original qubits by applying CNOT gates. Subsequently, multi-controlled rotation gates are employed to apply precise $R_{y}(\theta)$ operations to the auxiliary qubit, conditioned on the combined states of both the original and reversal qubits. For example, in the case of $n=2$, the proposed adaptive rotation circuit is illustrated in Fig. 7. When the original qubits are in $\left|10\right\rangle$, the reversal qubits are automatically in $\left|01\right\rangle$ due to the CNOT configuration. A controlled rotation gate is then activated based on this unique qubit configuration to apply the corresponding rotation $R_{y}(\theta_{2})$ to the auxiliary qubit. This implementation achieves high-precision realization of the controlled rotation operator $\mathcal{F}$. However, this accuracy comes at the cost of increased circuit depth and significant gate complexity, as $2^{n}$ controlled rotation gates are added to the circuit, which may hinder scalability for larger $n$.
Figure 7: Circuit-2 for precise implementation of $\mathcal{F}$.
@@
Step 3-Quantum Amplitude Estimation: Based on (12), we can apply the QAE algorithm to estimate the amplitude of the ancillary qubit being in $|1\rangle$, corresponding to the expectation of $\sum_{i=0}^{N-1}\hat{f}(i)p_{i}$.
@@
For brevity, we summarize the fundamental idea of QAE as follows: consider a quantum system composed of $n$ qubits in the state $|\psi\rangle_{n}$, along with an ancillary qubit initialized in the state $|0\rangle$. Let $\mathcal{A}$ denote a quantum operator acting on the qubits:
$\mathcal{A}|\psi\rangle_{n}|0\rangle=\sqrt{1-a}|\psi\rangle_{n}|0\rangle+\sqrt{a}|\psi\rangle_{n}|1\rangle.$ (13)
@@
QAE aims to estimate the probability of the “good” state $|1\rangle$ of the ancillary qubit, which is $a$. To achieve this goal, QAE employs the Grover operator $\mathcal{Q}=\mathcal{A}\mathcal{S}_{0}\mathcal{A}^{\dagger}\mathcal{S}_{\chi}$, where $\mathcal{S}_{0}$ is the operator for selective phase inversion of the initial state, and $\mathcal{S}_{\chi}$ is the operator for phase marking of the “good” state. The original amplitude estimation problem can be mapped to a phase estimation problem by iterative action on $\mathcal{Q}$. The amplitude $a\in[0,1]$ is encoded as a rotation angle $\theta\in[0,\frac{\pi}{2}]$, satisfying $a=\sin^{2}(\theta)$, and renders (13) as:
$\mathcal{A}|\psi\rangle_{n}|0\rangle=\cos(\theta)|\psi\rangle_{n}|0\rangle+\sin(\theta)|\psi\rangle_{n}|1\rangle.$ (14)
@@
By performing $k$ iterations of the Grover operator $\mathcal{Q}$, the state of the ancillary qubit is transformed to:
$\mathcal{Q}^{k}|\psi\rangle_{n}|0\rangle=\cos((2k+1)\theta)|\psi\rangle_{n}|0\rangle+\sin((2k+1)\theta)|\psi\rangle_{n}|1\rangle.$ (15)

### v2

@@ II-C DNs-ECs Coordinated Operation
Under limited information availability, DNs can only access the aggregated energy usage of each EC, while the detailed internal data and operational states of ECs remain unobservable. Therefore, the bilevel DNs-ECs coordinated operation problem cannot be solved directly. To address this challenge, we propose to learn direct mappings that capture the incentive-response behavior of each EC based on historical data, as in (2).
$\hat{P}^{\mathrm{EX,EC}}_{i}=\varphi_{i}(\pi^{\mathrm{EC}};\varTheta_{i})\sim\phi_{i}(\pi^{\mathrm{EC}}),$ (2)
@@
where the parameters $\varTheta_{i}$ can be optimized by minimizing a loss function (e.g., mean square error, MSE):
$\varTheta_{i}^{*}={\arg\min}_{\varTheta_{i}}\ \frac{1}{T}\sum_{t=1}^{T}\left(P^{\mathrm{EX,EC}}_{i,t}-\hat{P}^{\mathrm{EX,EC}}_{i,t}\right)^{2}.$ (3)
@@
Then, under the condition that the mapping model accurately captures the incentive-response behavior of ECs (i.e., the loss value is small), we can utilize the mapping model in (2) to approximate the unknown response function in (1h). The original bilevel DNs-ECs coordinated operation problem is reformulated as a single-level optimization problem, yielding solutions approximate to those of the original bilevel formulation.
By exploiting the differentiability of the mapping model (2) together with the CVaR-based formulation, in which auxiliary variables and small perturbations are introduced to convexify and eliminate the non-differentiability caused by the $\cdot$$_{+}$ operator, the integrated distribution network optimization model becomes fully differentiable and can therefore be effectively solved using gradient-based methods. The gradient descent update rule is:
$\pi^{\mathrm{EC},k+1}=\pi^{\mathrm{EC},k}-\eta\frac{\partial C}{\partial\pi^{\mathrm{EC},k}},\ \ v_{\alpha}^{k+1}=v_{\alpha}^{k}-\eta\frac{\partial C}{\partial v_{\alpha}^{k}},$ (4a)
$\frac{\partial C}{\partial\pi^{\mathrm{EC},k}}=\sum_{s=1}^{S}\gamma_{s}\frac{\partial C_{s}}{\partial\pi^{\mathrm{EC},k}}+\frac{\lambda}{1-\alpha}\sum_{s=1}^{S}\gamma_{s}\frac{\partial C_{s}}{\partial\pi^{\mathrm{EC},k}}\mathbb{I}(C_{s}>v_{\alpha}^{k}),$ (4b)
@@
$\frac{\partial C}{\partial v_{\alpha}^{k}}=\lambda\left(1-\frac{1}{1-\alpha}\sum_{s=1}^{S}\gamma_{s}\mathbb{I}(C_{s}>v_{\alpha}^{k})\right),$ (4c)
where $\mathbb{I}(C_{s}>v_{\alpha})=1$ if $C_{s}>v_{\alpha}$ and 0 otherwise. The feasible region of $\pi^{\mathrm{EC}}$ is ensured by repeatedly doing: firstly normalized to [0,1] via softmax then inverse-mapped to actual range to satisfy range constraint (1i) and mean-shifted to target mean to satisfy mean constraint (1j).
@@
    The response behavior of ECs should be accurately estimated in (2). However, this task is a complex long-horizon temporal sequence modeling problem, which poses significant challenges in both accuracy and generalizability.
@@
    Given $\pi^{\mathrm{EC},k}$, $\partial C_{s}/\partial\pi^{\mathrm{EC},k}$ has a closed form regarding the scenario value, but the gradient computation in (4) requires numerous scenarios to achieve accurate risk quantification, which leads to a heavy computational burden.
In the next section, inspired by recent advances in quantum computing, we propose a quantum learning and estimation approach to address the above two challenges, thus facilitating efficient DNs-ECs coordinated operation.
@@ III Methodology of
In this section, we present a quantum learning and estimation approach. Specifically, a Q-TCN-LSTM model is proposed to learn the incentive-response characteristics of ECs. Subsequently, a quantum estimation method is developed to accelerate the gradient computation across numerous scenarios.
Note that this study focuses on exploring the theoretical advantages of quantum learning and estimation. Given that current NISQ-era quantum hardware is constrained by noise, decoherence, and limited accessibility, practical deployment of the proposed methods is not yet feasible. Therefore, the subsequent analysis is performed on a simulated ideal quantum device, where external disturbances are excluded and hardware-level improvements (e.g., quantum error correction) are not considered.
III-A Quantum Learning
@@ III-A1 Variational Quantum Circuits
where $\left|\psi_{0}\right\rangle_{n}$ represents the initial quantum state of $n$ qubits; $\left|i\right\rangle_{n}$ is the $i$-th basis state in the $2^{n}$-dimensional Hilbert space, associated with amplitude $\alpha_{i}$ and a measurement probability of $\alpha_{i}^{2}$, which satisfies $\sum_{i=0}^{2^{n}-1}\alpha_{i}^{2}=1$. At this stage, VQC encodes exponentially many states via quantum superposition, offering computational advantages in high-dimensional problems. Besides, quantum entanglement enables a single quantum operation to act on all $2^{n}$ superposition states, thus significantly improving computational efficiency.
Secondly, the parameterized variational circuit block $V(\boldsymbol{\theta})$ applies a sequence of complex unitary transformations on $\left|\psi_{0}\right\rangle_{n}$. Single-qubit gates (e.g., Pauli-X gate, rotation gate) and multi-qubit gates (e.g., controlled-NOT gate, controlled-rotation gate) are used to manipulate the quantum state, collectively inducing quantum entanglement and further exploiting the superposition across qubits. This process enables the circuit to represent intricate nonlinear dependencies inherent in the input data. Usually, the variational circuit is composed of $L$ layers, each consisting of a set of parameterized or fixed single-qubit gates and entangling gates, forming the overall unitary:
@@
where $V^{\mathrm{ent}}_{(l)}$ is the entangling gate at layer $l$, and $V^{\mathrm{rot}}_{(l)}(\boldsymbol{\theta}_{(l)})=\bigotimes_{i=1}^{n}R_{(x,y,z)}(\theta_{(l),i})$ is the rotation gate at layer $l$. With the variational circuit, the quantum state is transformed as:
@@
where $\left\langle\psi(\boldsymbol{\theta})\right|$ is the conjugate transpose of $\left|\psi(\boldsymbol{\theta})\right\rangle$. Equation (8) denotes the expectation value of the observable Hermitian operator $\mathcal{O}$, resulting in classical scalar values for subsequent analysis.
The measurement output $\left\langle\mathcal{O}\right\rangle_{\boldsymbol{\theta}}$ represents a highly nonlinear transformation of the input, implicitly capturing complex correlations encoded by the VQC.
Regarding parameters optimization, classical optimization algorithms (e.g., Adam) are employed to compute the gradients of the predefined loss function (e.g., MSE for regression tasks) with respect to the trainable parameters, where gradients are approximated via methods such as the Parameter-Shift Rule. Parameters are iteratively updated to minimize the loss until convergence.
@@ III-A2 Structure of Q-TCN-LSTM
In response to time-varying electricity price signals, ECs exhibit both short-term (spanning several hours) response characteristics and long-term (over the scheduling horizon) globally optimal decision-making characteristics. Therefore, a short and long term coupling relationship exists between the response of ECs and price incentives. To capture this coupling, we propose a quantum-enhanced hybrid architecture, Q-TCN-LSTM, which integrates the advantages of temporal convolutional networks and long short-term memory networks. Specifically, TCNs can efficiently extract short-term features from time series data through causal and dilated convolution mechanisms, and can adjust the kernel size to control the receptive field. Then, LSTM can store long-term information in its cell states and selectively remember or forget information through gating mechanisms, effectively capturing long-term dependencies in time series. In the Q-TCN-LSTM model, the convolutional window of Q-TCN slides along the input time series, continuously extracting short-term coupling information within the sequence. Meanwhile, the original long sequence is transformed into a shorter sequence with a more compact dimension, which is then fed as input to the subsequent Q-LSTM layer to enable in-depth mining of data features. Fully connected (FC) layers then map the extracted features to the final output. The overall structure of the proposed Q-TCN-LSTM model is shown in Fig. 3. Notably, the structure of Q-TCN-LSTM (e.g., number of qubits, VQC depth, number of feature extraction layers) can be flexibly adjusted for different tasks.
@@ III-A3 Q-TCN for Short-Term Temporal Feature Extraction
We integrate VQCs into the causal convolutions to construct a Q-TCN model. For instance, a 4-qubit VQC for Q-TCN is illustrated in Fig. 4.
@@
Q-TCN utilizes a temporal rotational encoding scheme, where each input value $x_{t}$ is mapped to the angle of a rotation gate. This encoding preserves local temporal patterns while mitigating potential information loss commonly associated with sparse sampling in classical TCNs. Then, a parameterized variational circuit with skip-connected entanglement gates performs entanglement across temporally aligned qubits, constructing temporal dependencies without deep stacking or activations in classical TCNs. Finally, quantum measurements produce output features that capture interactions across multiple time steps. Quantum properties such as superposition and entanglement enhance Q-TCN’s expressive power with significantly fewer trainable parameters, effectively reducing model redundancy.
III-A4 Q-LSTM for Long-Term Temporal Feature Extraction
Based on the classical LSTM architecture, we construct a Q-LSTM model by replacing the gate mechanisms with VQCs, as illustrated in Fig. 5.
Figure 5: Structure of a Q-LSTM cell.
@@ III-A4 Q-LSTM for Long-Term Temporal Feature Extraction
where $\sigma$ and $\tanh$ blocks represent the sigmoid and the hyperbolic tangent activation functions, respectively. $x_{t}/h_{t}/c_{t}/y_{t}$ are the input, hidden state, cell state, and output at time $t$. $\otimes$ represents element-wise multiplication. $\cdot$] represents vector concatenation. $f_{t}$, $i_{t}$, $o_{t}$ represent the forget gate, input gate, and output gate in the LSTM model, respectively.
III-B Quantum Estimation
In this part, we first present the general methodology of quantum amplitude estimation. Then in light of the specific problem of this paper, we provide the detailed application scheme of QAE:
III-B1 General Methodology of QAE
Classical approaches typically rely on MC simulation for uncertainty estimation, which has been verified to require $\mathcal{O}(1/\varepsilon^{2})$ samples to achieve an estimation error of $\varepsilon$. This leads to substantial computational overhead. Alternatively, quantum amplitude estimation, leveraging quantum entanglement and amplitude amplification, can theoretically achieve the same precision level $\varepsilon$ with only $\mathcal{O}(1/\varepsilon)$ queries [29] under certain conditions [^1] Such conditions include the efficient preparation of the initial quantum state, small hardware noise in real devices, and stringent precision requirements, among others., offering a theoretical quadratic speedup compared to MC.
The QAE procedure generally involves three steps:
Step 1-Quantum State Preparation: A distribution $\mathcal{X}$ is divided into $N=2^{n}$ intervals, and is embedded into a quantum state:
@@ III-B1 General Methodology of QAE
where $|i\rangle_{n}$ represents a basis state of $n$ qubits, $p_{i}$ is the probability of state $|i\rangle_{n}$ with $\sum_{i=0}^{N-1}p_{i}=1$. The essence of preparing a quantum state embedded with distribution $\mathcal{X}$ lies in mapping probabilities to a quantum superposition state via amplitude encoding. In the uniform case where $p_{i}=1/N\ \forall i\in[0,\dots,N-1]$, (10) can be efficiently prepared using Hadamard gates. In more complex cases, it is often necessary to design sophisticated quantum state preparation circuits: for instance, constructing continuous Gaussian distributions using quantum arithmetic circuits [30, 31], or approximating complex distributions through parameterized quantum circuits [32].
Step 2-Quantum State Rotation: Assuming a function $f:\{0,1,...,2^{n}-1\}\rightarrow[0,1]$, a controlled rotation operator $\mathcal{F}$ is constructed such that
$\mathcal{F}|i\rangle_{n}|0\rangle=\sqrt{1-f(i)}|i\rangle_{n}|0\rangle+\sqrt{f(i)}|i\rangle_{n}|1\rangle.$ (11)
@@
$\mathcal{F}|\psi\rangle_{n}|0\rangle=\sum_{i=0}^{N-1}\sqrt{(1-f(i))p_{i}}|i\rangle_{n}|0\rangle+\sum_{i=0}^{N-1}\sqrt{f(i)p_{i}}|i\rangle_{n}|1\rangle.$ (12)
@@
In the implementation of $\mathcal{F}$, the auxiliary qubit undergoes a rotation $R_{y}(\theta)$ with $\theta=2\arcsin(\sqrt{f(i)})$, which means that $\theta$ is a nonlinear function of state $i$. Depending on the trade-off between circuit complexity and estimation accuracy, we propose two potential implementations of $\mathcal{F}$.
@@
Circuit-1 employs analytical approximations of the nonlinear rotation, as illustrated in Fig. 6. The core idea is to find a linear formulation with respect to $i$ that efficiently approximates $\arcsin(\sqrt{f(i)})$ such that $\theta\approx 2(ai+b)$. For example, when $x$ is sufficiently small, we have $\arcsin(\sqrt{x})\approx ax+b$, where $a$ and $b$ are determined using the least-squares method. Under this condition, the controlled rotation operator $\mathcal{F}$ can be implemented using a sequence of parameterized controlled-RY gates, with low implementation complexity. This approximation benefits from easy implementation but may introduce certain approximation errors.
Figure 6: The approximated implementation of $\mathcal{F}$, referred to as Circuit-1.
@@
Circuit-2 computes the exact nonlinear rotation angle corresponding to each quantum basis state. To this end, we design a rotation-angle-adaptive circuit architecture. In this scheme, $n$ original qubits are initialized to $|0\rangle_{n}$, $n$ reversal qubits to $|1\rangle_{n}$, and an auxiliary qubit to $|0\rangle$. The reversal qubits are kept in bitwise negated states of the original qubits by applying CNOT gates. Subsequently, multi-controlled rotation gates are employed to apply precise $R_{y}(\theta)$ operations to the auxiliary qubit, conditioned on the combined states of both the original and reversal qubits. For example, in the case of $n=2$, the proposed adaptive rotation circuit is illustrated in Fig. 7. When the original qubits are in $\left|10\right\rangle$, the reversal qubits are automatically in $\left|01\right\rangle$ due to the CNOT configuration. A controlled rotation gate is then activated based on this unique qubit configuration to apply the corresponding rotation $R_{y}(\theta_{2})$ to the auxiliary qubit. This implementation achieves high-precision realization of the controlled rotation operator $\mathcal{F}$. However, this accuracy comes at the cost of increased circuit depth and significant gate complexity, as $2n+1$ qubits are required and $2^{n}$ controlled rotation gates are added to the circuit, which may hinder scalability for larger $n$.
Figure 7: The precise implementation of $\mathcal{F}$, referred to as Circuit-2.
@@
Step 3-Quantum Amplitude Estimation: Based on (12), we can apply the QAE algorithm to estimate the amplitude of the ancillary qubit being in $|1\rangle$, corresponding to the expectation of $\sum_{i=0}^{N-1}f(i)p_{i}$. Commonly adopted QAE algorithms include canonical QAE, iterative QAE and maximum likelihood QAE. The percentage estimation error relative to the ground truth is defined as $\epsilon=$ $|estimation-truth|/truth\times 100\%$.
@@
For brevity, we summarize the fundamental idea of QAE [29] as follows. Consider a quantum system with $n$ qubits initialized in $|0\rangle_{n}$, and an ancillary qubit initialized in the state $|0\rangle$. Let $\mathcal{A}$ be a quantum unitary operator that:
$\mathcal{A}|0\rangle_{n}|0\rangle=\sqrt{1-a}|\psi_{0}\rangle_{n}|0\rangle+\sqrt{a}|\psi_{1}\rangle_{n}|1\rangle.$ (13)
@@
where $a\in[0,1]$ is unknown. $|\psi_{0}\rangle_{n}$ and $|\psi_{1}\rangle_{n}$ are two normalized states. QAE aims to estimate $a$ with high probability. To achieve this goal, QAE employs the Grover operator $\mathcal{Q}=\mathcal{A}\mathcal{S}_{0}\mathcal{A}^{\dagger}\mathcal{S}_{\mathcal{X}}$, where $\mathcal{S}_{0}$ is the operator for selective phase inversion of the initial state, and $\mathcal{S}_{\mathcal{X}}$ is the operator for phase marking of the “good” state. The original amplitude estimation problem can be mapped to a phase estimation problem by iterative action on $\mathcal{Q}$. The amplitude $a\in[0,1]$ is encoded as a rotation angle $\theta\in[0,\frac{\pi}{2}]$, satisfying $a=\sin^{2}(\theta)$, and renders (13) as:
$\mathcal{A}|0\rangle_{n}|0\rangle=\cos(\theta)|\psi_{0}\rangle_{n}|0\rangle+\sin(\theta)|\psi_{1}\rangle_{n}|1\rangle.$ (14)
@@
By performing $k$ iterations of the Grover operator $\mathcal{Q}$, the quantum state is transformed to:
$\mathcal{Q}^{k}\mathcal{A}|0\rangle_{n}|0\rangle=\cos((2k+1)\theta)|\psi_{0}\rangle_{n}|0\rangle+\sin((2k+1)\theta)|\psi_{1}\rangle_{n}|1\rangle.$ (15)